# LLM4AD: Large Language Models for Autonomous Driving — Concept, Review, Benchmark, Experiments, and Future Trends

### v4

@@ Abstract
With the broader adoption and highly successful development of Large Language Models (LLMs), there has been growing interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning capabilities, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to interactive decision-making. In this paper, we first introduce the novel concept of designing Large Language Models for Autonomous Driving (LLM4AD), followed by a review of existing LLM4AD studies. Then, we propose a comprehensive benchmark for evaluating the instruction-following and reasoning abilities of LLM4AD systems, which includes LaMPilot-Bench, CARLA Leaderboard 1.0 Benchmark in simulation and NuPlanQA for multi-view visual question answering. Furthermore, we conduct extensive real-world experiments on autonomous vehicle platforms, examining both on-cloud and on-edge LLM deployment for personalized decision-making and motion control. Next, we explore the future trends of integrating language diffusion models into autonomous driving, exemplified by the proposed ViLaD (Vision-Language Diffusion) framework. Finally, we discuss the main challenges of LLM4AD, including latency, deployment, security and privacy, safety, trust and transparency, and personalization.
@@ I Introduction
Large Language Models for Autonomous Driving
Concept of LLM4AD Review of Existing LLM4AD Studies Learning Method Pipeline Evaluation Deployment Platform Benchmark and Simulation LaMPilot-Bench CARLA Leaderboard 1.0 Benchmark NuplanQA Benchmark Summary Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making Talk2Drive Framework Autonomous Driving System Experiment Setup Experiment Results Real-Vehicle Experiments: On-Edge VLM for Personalized Motion Control Problem Statement Our Framework Real-World Experiment Future Trend of LLM4AD: Language Diffusion Models in Autonomous Driving ViLaD: Vision Language Diffusion Experiment and Results Real-World Case Studies Challenges Latency Security and Privacy Development Safety Personalization Trust and Transparency
Figure 1: A condensed overview of this paper.
@@ I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from the classic natural language processing (NLP) tasks like document modification and information extraction to the new emerging scenarios such as LLM-based agents and LLMs for evaluation [zhao2023survey]. One particular application that our study focuses on, like many other domains, is adopting LLMs for autonomous driving (LLM4AD). In this area, various advanced emerging LLM-based algorithms and technologies are continually enhancing the capabilities of autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency forward. LLMs can contribute to autonomous systems from high-level decision-making processes to meticulous low-level control.
@@
On the high-level side, LLMs can actively engage in adjusting driving mode or decision-making process [cui_drive_2024]. Imagine the scenario in which you are seated in an autonomous vehicle, where all you need to do is to abstractly express your feelings like “I do not want my friends to wait for me.” The LLM4AD system then interprets your emotions and adjusts the vehicle’s control strategies accordingly, aligning with your current driving mood or preferences. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret humans’ intentions only from some vague expressions [cui2023large]. Besides, LLM-based systems have significant potential for achieving personalization in autonomous driving due to their continuous learning capability. This capability enables it to continuously adapt to individuals’ preferences and desires, improving the driving experience according to different users. In addition, LLMs also hold the potential to develop a knowledge-driven system capable of driving like an expert, which can accumulate experience through continuous driving [fu_drive_2024].
@@
On the other hand, on the low-level side, LLMs can also play a crucial role in the tuning and control process. LLMs have demonstrated the ability to analyze specific scenarios and convert gathered information into mathematical representations to guide low-level controllers [sha_languagempc_2023]. Additionally, LLMs can receive input data from the controller and provide performance updates to aid humans in analyzing the control loop’s effectiveness [s23135899]. They can potentially suggest improvements or detect issues to enhance overall performance.
@@
Besides applying LLMs into the conventional module-based autonomous systems, LLMs also play a significant role in the end-to-end system for autonomous driving. By leveraging their powerful reasoning and multimodal understanding capabilities, LLMs can directly process sensory data and ego-vehicle states to generate planning trajectories and fine-grained control actions [chen2024end]. This LLM-driven paradigm has attracted increasing attention from both industry [driveirl, hu2023imitation, pmlr-v164-scheel22a] and academia [hu_planning-oriented_2023]. Unlike the conventional modular-based paradigms that separately handle perception, prediction, planning, and control, the LLM-based end-to-end autonomous driving system is fully differentiable and optimizes the entire driving process simultaneously.
@@
The emerging developments of LLM4AD systems raise a crucial question for all related researchers: Why have LLMs become incredibly popular in autonomous driving? What are the advantages of LLM4AD systems over those without LLM integration? Based on current research and developments, we summarize several compelling advantages of integrating LLMs as follows:
@@
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Humans can express abstract commands and feelings to the vehicle, while the LLMs can accurately capture the intentions behind human expressions.
@@
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources such as traffic rules and accident reports, allowing generated decisions to guarantee safety and follow local regulations.
@@
  • Zero-Shot and Few-Shot Planning: The zero-shot generalization ability enables LLMs to perform tasks they have not been trained on before [kojima_large_2022]. This means LLM-based autonomous systems can handle uncommon situations with minimal or no prior experience, enabling them to confidently navigate through “corner cases.”
@@
  • Continuous Learning and Personalization: LLMs learn and adapt continuously, providing the ability to follow individual user preferences and improve the driving experience over time.
@@
  • Interpretability and Trust: LLMs can explain their decisions in natural language, improving trust and understanding between autonomous systems and their users.
@@
Despite the numerous advantages above, LLMs still have some limitations and potential issues. The most obvious one is sometimes they are inapplicable to real-time tasks in autonomous driving. Specifically, LLMs typically require seconds to process textual information. This delay can pose significant safety concerns and risks in situations requiring immediate, real-time decision-making. Additionally, how to avoid the negative effect provided by hallucinations is also a main challenge in LLM4AD systems. “Hallucination” refers to instances where LLMs generate output that is factually incorrect, nonsensical, or unrelated to the input prompt. As autonomous driving is a safety-critical task, hallucinations can introduce significant safety and reliability concerns and potentially undermine trust in LLM4AD systems. Furthermore, data privacy and security are other potential issues since LLMs collect and process vast amounts of textual data, including potentially sensitive information about their surroundings, passengers, and driving preferences.
@@ I-B Overview of This Article
Fig. 1 provides an overview of the structure of this paper. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III reviews existing works related to the integration of LLMs into autonomous driving systems. Sec. IV develops open benchmarks designed to evaluate the performance of LLMs in various autonomous driving tasks, which includes datasets, evaluation metrics, and baseline models. Sec. V and Sec. VI present the implementation of our framework on a real-vehicle platform and discuss the results of a series of experiments conducted to validate the system’s driving and personalization performance. Sec. VII proposes the future trend of leveraging large language diffusion models in autonomous driving. Sec. VIII analyzes the current challenges and outlines promising research directions for leveraging LLMs in autonomous driving applications. Finally, Sec. IX summarizes the key contributions and findings of this work.
Figure 2: The conceptual illustration of the proposed LLM4AD system.
@@ II Concept of LLM4AD
In this section, we introduce a perspective where LLMs play the role of the decision-making “brain” within the autonomous driving system. Within our framework, LLMs do not directly affect perception or localization modules, which serve as the vehicle’s “eyes;” instead, they utilize outputs from these modules as references to guide the high-level decision-making process. By receiving processed data from them, LLMs can enhance informed decision-making, leading to significant improvements in the performance of autonomous vehicles. At the downstream, the vehicle’s control module serves as its “hands,” executing the driving policy obtained from the LLM-based decision-making process.
The overall LLM4AD framework is shown in Fig. 2. In our framework, humans provide instructions and evaluations, where the instructions $I$ and feedback/evaluations $F$ along with historical memories $H$, system messages $S$, and context information $C$ serve as inputs to LLMs. The memory module stores corresponding historical interactions between humans and vehicles $H$ for different users. Upon receiving these inputs, LLMs engage in reasoning and produce outputs, including generated Language Model Programs (LMPs) or policies $P$ and reasoning thoughts $R$. The $P$ is then sent to the executor for execution in the environment, while the reasoning thoughts help LLMs generate more reasonable driving policies. It is important to note that this is a general concept, where specific implementations may vary across different applications. However, all applications follow the principles and framework of this concept.
@@ II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are directly fed into LLMs in natural language. $I$ includes the humans’ desired needs for the autonomous agent, while human evaluation $F$ is the feedback on the effectiveness of these driving policies.
@@ II-B System Message
The system messages $S$ provide instructions or context to the LLM4AD system at the start of a conversation or task. System messages $S$ allow developers to clearly define the role, constraints, and objectives that LLM4AD systems should follow when engaging with users or working on particular types of tasks.
System messages $S$ in autonomous driving tasks act like a set of high-level guidelines or rules that frame how LLMs should operate and make decisions. These high-level system messages are crucial in defining the basic driving logic, including tasks such as task definition, adherence to traffic rules, description of decision states, and overall goals or metrics to optimize for. They serve as a foundational framework guiding the behavior and decision-making processes of autonomous vehicles on the road. Without a well-crafted system message, LLMs could make incorrect assumptions or unintended strategies.
@@ II-C Situation Descriptor
A situation descriptor transfers the current driving context $C$ into textual descriptions. The situation descriptor follows a predefined structure and employs natural language to provide a detailed textual description of the current driving scenario. It aims to provide LLMs with situational awareness and a comprehensive representation of the current driving scenario, allowing them to make appropriate decisions in the current traffic situation. This could include descriptive statements like “You are positioned in the leftmost lane of a two-way highway” or “A vehicle is located 50 meters ahead of your current position.” The descriptor translates complex spatial and temporal relationships between different road users and ego vehicles intuitively into the natural language format, allowing context information to be reasoned by LLMs.
@@ II-D History Memory and Memory Module
The memory module stores profiles from various users to enhance driving experiences for all users. Whenever a human user utilizes the LLM4AD system, the system logs the relevant historical interactions $H$ associated with that user. Subsequently, the historical data for the current human user is transmitted to LLMs as input. This particular historical interaction $H$ serves as a reference point for the preferences of the current users, therefore guiding the system to improve user experience. After each trip, the interaction data will be updated in the corresponding profile in the memory module. The specifics of the interaction data vary for each system and will be elaborated in their respective sections.
@@ II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human feedback $F$ and generate the textual outputs (Generated Policies/LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
@@ II-F Generated Policy/Program
Inspired by the concept of “Code as Policy [liang_code_2023],” one of the key outputs from LLMs is the generated Policies/s $P$ consisting of executable codes. These codes are used to affect the driving behavior of the ego agent within the environment, and they not only have the ability to generalize to new natural language commands but can also provide precise numerical values such as velocities based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
@@ II-G Output Thoughts
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is ‘hurry,’ I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated policies/LMPs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.
@@ II-H Executor
The executor acts as the bridge between LLMs’ textual outputs and the current autonomous driving policy. It takes the generated policies/LMPs $P$ from LLMs and executes it in the corresponding environment. This allows the codes to interact with the current status of the ego vehicle and enables the $P$ to deploy their intended driving behaviors in the real or simulated environment. In our scope, different autonomous driving systems will have some differences in how those codes are executed by their respective executors.
@@ III Review of Existing LLM4AD Studies
This section reviews existing studies that apply LLMs for autonomous driving, which are grouped into four aspects: learning method, pipeline design, evaluation methodology, and deployment platform. On the learning side, recent work spans supervised fine-tuning with parameter-efficient variants, prompt engineering strategies, retrieval-augmented generation, imitation learning, and reinforcement learning from human or environment feedback, reflecting a progression from static adaptation to closed-loop policy optimization. The pipeline discussion highlights how these frameworks select and combine input modalities (textual scene descriptions, camera images, LiDAR point clouds, and vectorized driving states) and map them to diverse output representations, including natural language explanations, discrete meta-actions, low-level control signals, trajectories, and even policy code. For evaluation, existing studies rely on language-augmented driving datasets, vision-language benchmarks, simulation environments, and a small but growing set of real-vehicle deployments, using both task-specific driving metrics and natural-language quality metrics, yet still lack a unified benchmark for LLM4AD. Finally, we summarize real-world deployment platforms and compare their capabilities, showing that most current testbeds specialize in specific autonomy components, while only very recent efforts begin to support full-stack validation tailored to LLM-integrated autonomous driving.

### v5

@@ Abstract
With the broader adoption and highly successful development of Large Language Models (LLMs), there has been growing interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning capabilities, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to interactive decision-making. This paper first introduces the novel concept of designing Large Language Models for Autonomous Driving (LLM4AD), followed by a review of existing LLM4AD studies. Then, a comprehensive benchmark is proposed for evaluating the instruction-following and reasoning abilities of LLM4AD systems, which includes LaMPilot-Bench, CARLA Leaderboard 1.0 Benchmark in simulation and NuPlanQA for multi-view visual question answering. Furthermore, extensive real-world experiments are conducted on autonomous vehicle platforms, examining both on-cloud and on-edge LLM deployment for personalized decision-making and motion control. Next, the future trends of integrating language diffusion models into autonomous driving are explored, exemplified by the proposed ViLaD (Vision-Language Diffusion) framework. Finally, the main challenges of LLM4AD are discussed, including latency, deployment, security and privacy, safety, trust and transparency, and personalization.
@@ I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from classic natural language processing (NLP) tasks, such as document modification and information extraction, to emerging scenarios like LLM-based agents and evaluation frameworks [zhao2023survey].
A prominent application is the adoption of LLMs for autonomous driving (LLM4AD). In this area, various advanced algorithms are continually enhancing autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency. These models can contribute to autonomous systems ranging from high-level decision-making processes to meticulous low-level control.
@@
On the high-level side, LLMs can actively engage in adjusting driving modes or decision-making process [cui_drive_2024]. Consider a scenario where a passenger abstractly expresses urgency, such as “I do not want my friends to wait for me.” An LLM4AD system can interpret these emotions and adjust the vehicle’s control strategies accordingly, aligning with the current driving mood. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret human intentions derived soly from vague expressions [cui2023large]. Furthermore, LLM-based systems offer significant potential for personalization due to their continuous learning capability. This enables the system to adapt to individual preferences, improving the driving experience for different users. Additionally, LLMs hold the potential to develop knowledge-driven systems capable of driving like experts by accumulating experience through continuous operation [fu_drive_2024].
@@
On the low-level side, LLMs play a crucial role in the tuning and control process. They have demonstrated the ability to analyze specific scenarios and convert gathered information into mathematical representations to guide low-level controllers [sha_languagempc_2023]. Additionally, LLMs can receive input data from the controller and provide performance updates to aid in analyzing the control loop’s effectiveness [s23135899], potentially suggesting improvements or detecting issues to enhance overall performance.
@@
Beyond applying LLMs to conventional modular autonomous systems, they also play a significant role in end-to-end systems. By leveraging powerful reasoning and multimodal understanding capabilities, LLMs can directly process sensory data and ego-vehicle states to generate planning trajectories and fine-grained control actions [chen2024end]. This LLM-driven paradigm has attracted increasing attention from both industry [driveirl, hu2023imitation, pmlr-v164-scheel22a] and academia [hu_planning-oriented_2023]. Unlike conventional modular paradigms that handle perception, prediction, planning, and control separately, the LLM-based end-to-end system is fully differentiable and optimizes the entire driving process simultaneously.
@@
The emerging developments in LLM4AD raise crucial questions: Why have LLMs become popular in autonomous driving, and what advantages do they offer over systems without LLM integration? Based on current research, compelling advantages of integrating LLMs include:
@@
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Users can express abstract commands and feelings, while the LLM accurately captures the intentions behind these expressions.
@@
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources, such as traffic rules and accident reports, allowing generated decisions to ensure safety and adherence to local regulations.
@@
  • Zero-Shot and Few-Shot Planning: Zero-shot generalization enables LLMs to perform tasks on which they have not been explicitly trained [kojima_large_2022]. Consequently, LLM-based systems can handle uncommon situations with minimal prior experience, enabling confident navigation through “corner cases.”
@@
  • Continuous Learning and Personalization: LLMs learn and adapt continuously, facilitating the ability to follow individual user preferences and improve the driving experience over time.
@@
  • Interpretability and Trust: LLMs can explain decisions in natural language, thereby improving trust and transparency between autonomous systems and users.
@@
Despite these advantages, LLMs possess limitations. A primary concern is their applicability to real-time tasks; LLMs typically require seconds to process textual information. This latency poses safety risks in situations demanding immediate decision-making. Additionally, preventing hallucinations, where LLMs generate factually incorrect or nonsensical outputs, remains a challenge. As autonomous driving is safety-critical, hallucinations can undermine system reliability and user trust. Furthermore, data privacy and security present issues, as LLMs collect and process vast amounts of textual data, including potentially sensitive information regarding surroundings, passengers, and driving preferences.
@@ I-B Overview of This Article
The remainder of this paper is organized as follows: Sec. II introduces the concept and key elements of the proposed LLM4AD framework. Sec. III reviews existing works related to the integration of LLMs into autonomous driving systems. Sec. IV establishes open benchmarks designed to evaluate the performance of LLMs in various tasks, including datasets, evaluation metrics, and baseline models. Sec. V and Sec. VI present the implementation of the framework on a real-vehicle platform, discussing the results of extensive experiments conducted to validate driving and personalization performance. Sec. VII explores future trends in leveraging large language diffusion models for autonomous driving. Sec. VIII analyzes current challenges and outlines promising research directions, and finally, Sec. IX summarizes the key contributions and findings of this work.
Figure 1: The conceptual illustration of the proposed LLM4AD system. The process begins with five distinct inputs being fed into the core large language models: historical memory ($H$) retrieved from a memory module, system messages ($S$) that define the role and constraints, a situation descriptor ($C$) providing context, human instructions (II), and human evaluations ($F$). These inputs are processed by models such as LLaMA, Gemini, or Claude to generate two primary outputs: reasoning thoughts ($R$) that explain the decision-making process and language model programs ($P$) which serve as executable policies. An executor then acts as a bridge, translating these programs into actions within the environment, which can range from simulators like CARLA and HighwayEnv to real-world autonomous driving platforms. Finally, the system incorporates a feedback loop where historical interactions are stored back into the memory module to support continuous learning and personalization.
@@ II Concept of LLM4AD
This section introduces a perspective in which LLMs function as the decision-making “brain” within the autonomous driving system. In the proposed framework, LLMs do not directly alter perception or localization modules (the vehicle’s “eyes”); instead, they utilize outputs from these modules as references to guide high-level decision-making. By processing this data, LLMs enhance informed decision-making, leading to significant performance improvements. Downstream, the vehicle’s control module serves as the “hands,” executing the driving policy derived from the LLM-based process.
The overall LLM4AD framework is illustrated in Fig. 1. Humans provide instructions $I$ and evaluations/feedback $F$. These, along with historical memories $H$, system messages $S$, and context information $C$, serve as inputs to the LLMs. The memory module stores the corresponding historical interactions $H$ for different users. Upon receiving these inputs, the LLMs engage in reasoning to produce outputs, including generated Language Model Programs (LMPs) or policies $P$ and reasoning thoughts $R$. The policy $P$ is sent to the executor for implementation in the environment, while the reasoning thoughts assist the LLMs in generating more logical driving strategies. While specific implementations may vary across applications, they all adhere to the principles of this general concept.
@@ II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are fed directly into the LLMs in natural language. $I$ represents the human’s desired objectives for the autonomous agent, while the evaluation $F$ provides feedback on the effectiveness of the executed driving policies.
@@ II-B System Message
System messages $S$ provide instructions and context to the LLM4AD system at the initiation of a conversation or task. They allow developers to clearly define the role, constraints, and objectives the system must follow. In autonomous driving tasks, $S$ acts as a set of high-level guidelines framing how LLMs should operate, defining basic driving logic such as task definitions, adherence to traffic rules, decision state descriptions, and optimization metrics. These messages serve as a foundational framework to prevent incorrect assumptions or unintended strategies.
@@ II-C Situation Descriptor
The situation descriptor translates the current driving context $C$ into structured natural language descriptions. Its objective is to equip LLMs with situational awareness and a comprehensive representation of the driving scenario — such as “You are positioned in the leftmost lane of a two-way highway” or “A vehicle is located 50 meters ahead” — enabling appropriate decision-making. This descriptor intuitively translates complex spatial and temporal relationships between road users and the ego vehicle into a format the LLM can reason about.
@@ II-D History Memory and Memory Module
The memory module stores user profiles to enhance the driving experience. Whenever a human utilizes the LLM4AD system, relevant historical interactions $H$ associated with that user are logged. Subsequently, this historical data is transmitted to the LLMs as input, serving as a reference point for user preferences and guiding the system to improve the user experience. Interaction data is updated in the corresponding profile within the memory module after each trip.
@@ II-E Large Language Models
LLMs serve as the core module of the framework, receiving all inputs (history interactions $H$, system message $S$, situation description $C$, human instruction $I$, and human feedback $F$) to generate textual outputs (Generated Policies/LMPs $P$ and reasoning thoughts $R$). Notably, chain-of-thought prompting [wei_chain--thought_2022] is employed as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
@@ II-F Generated Policy/Program
Inspired by the concept of “Code as Policy” [liang_code_2023], a primary output of the LLMs is the generated Policy $P$, consisting of executable code. These codes influence the ego agent’s driving behavior within the environment. They possess the ability to generalize to new natural language commands and provide precise numerical values (e.g., velocities) based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
@@ II-G Output Thoughts
Through chain-of-thought prompting, LLMs generate not only the program code but also step-by-step explanations of the thought process behind the solution. These chains of thought $R$ elucidate the reasoning behind decisions (e.g., “Since the command is ‘hurry,’ I will increase the target velocity”). The output thoughts $R$ accompany the generated policies $P$, offering insights into how natural language commands are interpreted within the specific driving context to produce precise control values, thereby improving system transparency and interpretability.
@@ II-H Executor
The executor bridges the gap between the LLMs’ textual outputs and the autonomous driving policy. It receives the generated policies/LMPs $P$ and executes them in the corresponding environment, allowing the code to interact with the ego vehicle’s status and deploy the intended behaviors in real or simulated settings. Execution specifics may vary depending on the underlying autonomous driving system.
@@ III Review of Existing LLM4AD Studies
This section reviews existing studies that apply LLMs for autonomous driving, which are grouped into four aspects: learning method, pipeline design, evaluation methodology, and deployment platform. On the learning side, recent work spans supervised fine-tuning with parameter-efficient variants, prompt engineering strategies, retrieval-augmented generation, imitation learning, and reinforcement learning from human or environment feedback, reflecting a progression from static adaptation to closed-loop policy optimization. The pipeline discussion highlights how these frameworks select and combine input modalities (textual scene descriptions, camera images, LiDAR point clouds, and vectorized driving states) and map them to diverse output representations, including natural language explanations, discrete meta-actions, low-level control signals, trajectories, and even policy code. For evaluation, existing studies rely on language-augmented driving datasets, vision-language benchmarks, simulation environments, and a small but growing set of real-vehicle deployments, using both task-specific driving metrics and natural-language quality metrics, yet still lack a unified benchmark for LLM4AD. Finally, this section summarizes real-world deployment platforms and compares their capabilities, showing that most current testbeds specialize in specific autonomy components, while only very recent efforts begin to support full-stack validation tailored to LLM-integrated autonomous driving.